OpenStudy (anonymous):

(x-5/2)^2 = 15/4 Use SQUARE ROOT PROPERTY

OpenStudy (anonymous):

\[(x-5/2)^2 = 15/4 \implies x-5/2 = \pm \frac{\sqrt{15}}{2}\]

OpenStudy (anonymous):

So final answer is x= 5 +/- (sq.root) 15 ----------- 2 ?

OpenStudy (anonymous):

Maybe